OpenStudy (anonymous):

If you toss two dice 1,000 times, how many times do you expect to have the sum of the two dice equal to four? about 250 about 167 about 83 about 42

OpenStudy (blacksteel):

Each die has 6 different possible values it can show. Then if you roll two dice, there are 6*6 = 36 possible combinations of numbers you can get. Of these 36 combinations, there are 3 ways to roll a 4: 1 on the first die, 3 on the second 2 on both dice 3 on the first die, 1 on the second Thus, you have a 3/36 = 1/12 chance of rolling a total value of 4. Thus, if you roll the dice 1000 times, you expect to get a sum of 4 1000/12 = 83 1/3 ~ 83 times.